A ray of light travels from air into another mediu
Answer #1

Let Refractive index =\mu

\mu =\frac{\sin \theta _{1}}{\sin \theta _{2}}

\theta _{1}= Angle of incidence =45^{0}

\theta _{2}= Angle of refraction

(a)

\mu =2.419

\sin \theta _{2}=\frac{\sin 45^{0}}{2.419}=0.2923

\theta _{2}=\sin^{-1}(0.2923)=17^{0}

(b)

\mu =1.628

\sin \theta _{2}=\frac{\sin 45^{0}}{1.628}=0.4343

θ2 sin-1 (0.4343)-5.740

(c)

\mu =1.544

\sin \theta _{2}=\frac{\sin 45^{0}}{1.544}=0.458

θ2 sin-1 (0.458-27.260